MIR EXTENSION OPERATIONS WEEKLY REPORT FOR WEEK OF Monday, 12.06.95 through Friday, 16.06.95
Report filed by Sarah Kirby
The crew spent the day at rest. They each enjoyed a private conference with their families. A few medical experiments and some minor EVA preparations were also conducted.
The crew burned three O2 candles today. Late in the day, the commands to activate the Kvant-2 array articulation drive were read up. The crew was told that the last time the drive was activated a unit started to overheat. The crew was asked to monitor that unit and if it looked like it was overheating to tell the ground because they had some plans to provide fans for more active cooling of the unit.
The comm session through Dryden did not work. Dryden reported a good uplink, but no downlink (not even static) was received so it is entirely possible that the crew chose not to participate. Later in the day Soloviev cancelled all American ground pass attempts until at least next Monday.
There is a problem in the packet transmission from both the ground and the crew. Therefore the crew reported on the open loop a problem that occurred during the last move of Kristall (the core module shook). The crew also reported that one of the spare suits (which they are evidently preparing for the EVA) "stinks". They had already been blowing air through it for two days.
Soloviev asked the crew to do their best to remove the TREK detector on the next EVA if time permits at the end. Later the crew said that the pressure was "up to the mark" and that they were feeling much better. Finally, the crew did a test on the Strela to see if they could tell whether it was hindering the Kvant-2 array. The crew said they thought that the boom might be hindering the array but that the EVA should provide a definite answer.